How much water will spill from a 5.00 L vat if the temperature increased by 35.0°C?

1
Expert's answer
2021-12-27T08:01:48-0500

The initial volume of the vat is "V_0." By increasing the temperature, the volume of the vat increases, decreasing the inner space:


"V=V_0(1+\alpha \Delta t)^3,\\
\Delta V=V-V_0=V_0\alpha^3 t^3,\\
\Delta V=5\alpha^3·35^3=214375\alpha ^3."

For a copper vat, the volume of the spilled water will be 9.98·10-10 L.